# Models, Methods and Waveforms for Estimation and Prediction of Doubly Sparse Time-Varying Channels

### v1

@@ Appendix B Proof of Theorem 2
where $\widetilde{\mathbf{y}}_{\rm p}=[\widetilde{\mathbf{y}}_{{\rm p},0}^{\rm T}\quad\cdots\quad\widetilde{\mathbf{y}}_{{\rm p},(L-1)P}^{\rm T}]^{\rm T}$. For each $l$, $\widetilde{\mathbf{y}}_{{\rm p},l}$ is a $N_{\rm p}\times 1$ vector composed of the pilot samples received at DAFT domain positions $\left\{m_{p}+l\right\}_{p=1\cdots N_{\rm p}}$. Note that by this definition $\tilde{\mathbf{y}}_{\rm p}$ is obtained by permuting $\mathbf{y}_{\rm p}$ in (29) in accordance with the permutation that gives $\tilde{\boldsymbol{\alpha}}$ from $\boldsymbol{\alpha}$. Next, by setting for each $p\in\llbracket 1\mkern 1.5mu..\mkern 1.5muN_{\rm p}\rrbracket$ $m_{p}=q_{p}\frac{N}{2\lceil\frac{Q}{P}\rceil+1}$ for some integer $q_{p}$ it follows from (14) and (43) that $\widetilde{\mathbf{M}}_{\rm p}$ has the following Kronecker structure
$\widetilde{\mathbf{M}}_{\rm p}=\mathbf{I}_{(L-1)P+1}\otimes\widetilde{\mathbf{M}}_{\cal D},$ (46)
with $\widetilde{\mathbf{M}}_{\mathcal{D}}=\mathrm{diag}({\rm p}_{1}\cdots{\rm p}_{N_{\rm p}})\mathbf{F}_{2\lceil\frac{Q}{P}\rceil+1,N_{\rm p}}\boldsymbol{\Psi}$, $\mathbf{F}_{2\lceil\frac{Q}{P}\rceil+1,N_{\rm p}}$ is a $\left(2\lceil\frac{Q}{P}\rceil+1\right)\times N_{\rm p}$ partial Fourier measurement matrix and $\boldsymbol{\Psi}$ is a diagonal matrix with unit-modulus entries. We can thus use [28, Theorem 4.5] pertaining to sub-sampled Fourier matrices to get that for sufficiently large $L$, $Q$, sufficiently small $\delta$, and
$N_{\rm p}>O\left(\frac{1}{\delta^{2}}\log^{2}\frac{1}{\delta}\log\frac{\log(LP)}{\delta}\log(LP)\log\frac{Q}{P}\right)$ (47)
the RIP constant $\delta_{\tilde{s}_{\rm D}}$ of $\widetilde{\mathbf{M}}_{\mathcal{D}}$ satisfies $\delta_{\tilde{s}_{\rm D}}\leq\delta$ with probability $1-e^{-\Omega\left(\log{\frac{Q}{P}}\log{\frac{1}{\delta}}\right)}$. The RIP of $\mathbf{I}_{(L-1)P+1}$ trivially satisfies $\delta_{\tilde{s}_{\rm d}}=0$. As for the HiRIP of $\widetilde{\mathbf{M}}_{\rm p}$, we can apply [18, Theorem 4] to (46) thanks to its Kronecker structure to get that, f $N_{\rm p}$ and $\delta$ are as in (47), then
@@ Appendix C Proof of Theorem 3
We first need the following lemma taken from [29, Theorem 2.4].
@@
[29, Theorem 2.4] Let $h(t)$ be a continuous-time zero-mean wide-sense stationary random process with power spectrum $P_{h}(f)=\frac{1}{B}\mathrm{rect}(\frac{f-F_{\rm c}}{2B})$. Denote by $\mathbf{h}=[h(0T_{\rm s})\,\ldots\>h((N-1)T_{s})]^{\rm T}$ a vector of samples acquired from $h(t)$ with a sampling period $T_{s}\leq\frac{1}{2F_{\rm c}+B}$. Let $W=\frac{BT_{s}}{2}$, $\mathbf{U}_{k}$ the matrix form of the $k$ first $(N,W)$-DPSS vectors, $\mathbf{E}_{f}\triangleq\mathrm{diag}\left(e^{\imath 2\pi f0}\ldots e^{\imath 2\pi f(N-1)}\right)$ and $\mathbf{P}_{k}\triangleq\mathbf{E}_{F_{\rm c}T_{\rm s}}\mathbf{U}_{k}\mathbf{U}_{k}^{\rm H}\mathbf{E}_{F_{\rm c}T_{\rm s}}^{\rm H}$. Then $\mathbb{E}\left[\left\|\mathbf{h}-\mathbf{P}_{k}\mathbf{h}\right\|_{2}^{2}\right]=\frac{1}{2W}\sum_{l=k}^{N-1}\lambda_{l}^{(N,W)}$.
In what follows, we use Lemma 2 to upper bound $\mathbb{E}\left[\left|h_{l,q,n}-h_{l,q,n}^{\rm BEM}\right|^{2}\right]$. For that sake, we rewrite $h_{l,q,n}$ as the sampled version of the continuous-time signal $h_{l,q}(t)$ defined as
@@
with $f_{i}\triangleq\frac{q}{NT_{\rm s}}+\frac{\kappa_{i}}{NT_{\rm s}}$. To prove that the PSD of the random process $h_{l,q}(t)$ has the desired property, we derive its autocorrelation function $R_{h_{l,q}}(\tau)\triangleq\mathbb{E}[h_{l,q}(t)h_{l,q}^{*}(t+\tau)]$ as
$R_{h_{l,q}}(\tau)$ $=\sum_{i=1}^{N_{D}}\sum_{j=1}^{N_{D}}\mathbb{E}[\alpha_{l,q,i}\alpha_{l,q,j}^{*}]\,\mathbb{E}\left[e^{\imath 2\pi f_{i}t}e^{-\imath 2\pi f_{j}(t+\tau)}\right]$
@@
where the first equality is due to (49) and the second to $f_{i}\sim\mathcal{U}\left(\left[\frac{q}{NT_{\rm s}}-\frac{1}{2NT_{\rm s}},\frac{q}{NT_{\rm s}}+\frac{1}{2NT_{\rm s}}\right]\right)$, $\alpha_{l,q,i}\sim\mathcal{CN}(0,\sigma_{\alpha}^{2})$ and the independence property of $\left\{\alpha_{l,q,i}\right\}_{i}$ as per Definition 4. This gives:
@@
The power spectral density (PSD) is thus
$P_{h_{l,q}}(f)\triangleq\mathcal{F}\{R_{h_{l,q}}(\tau)\}=N_{D}\,\sigma_{\alpha}^{2}\,NT_{\rm s}\,\text{rect}\left(\left(f-\frac{q}{NT_{\rm s}}\right)NT_{\rm s}\right).$ (52)
@@
$\mathbb{E}\left[\left|h_{l,q,n}-h_{l,q,n}^{\rm BEM}\right|^{2}\right]=\frac{1}{N}\mathbb{E}\left[\left\|\mathbf{h}_{l,q}-\mathbf{h}_{l,q}^{\rm BEM}\right\|^{2}\right]=\frac{N_{\rm D}\,\sigma_{\alpha}^{2}}{2WN}\sum_{b=Q_{\rm BEM}}^{N-1}\lambda_{b}^{(N,W)}\>.$ (53)
Now define $\lambda_{b}^{(c)}$ as the $b$-th eigenvalue of the prolate spheroidal wave functions (PSWF) [30] with the bandwidth parameter $c\triangleq\pi NW$. This allows us to exploit existing results on the behavior of PSWF eigenvalues in the limit of $c\to\frac{\pi}{2}$ (or equivalently as $W$ tends to zero at the rate $\frac{1}{2N}$) to upper bound the sum of DPSS eigenvalues $\lambda_{b}^{(N,W)}$ in (53). Indeed, due to [31, Theorem 2]
@@
Here, $A_{W}$ is a function of $W$ defined in [31, Eq. (45)] and whose image is fully included in the interval $\left[\frac{\pi^{2}}{8},2\right]$. Plugging (54) into (53) and noting that $12WN=$ 1 lead to
@@
The right-hand side term in (55) can be upper bounded due to the fact that the PSWF eigenvalues decay at least exponentially [^1] Actually, even super-geometric decay can be proven [32] as $b$ grows beyond $\frac{2\pi NW}{\pi}+O(\log(\pi NW))=1+O(\log\frac{\pi}{2})$. More precisely, it follows from [33, Theorem 2.5] that $\lambda_{b}^{(c)}=O\left(e^{-\frac{\pi^{2}}{\log\frac{\pi}{2}}b}\right)$. Plugging this into (55) results, for any $\epsilon>0$ and $Q_{\rm BEM}>C\log\frac{2}{\epsilon}$ for sufficiently large $C$, in
@@
$\mathbb{E}\left[\sum_{l=0}^{L-1}\left|h_{l,n}-h_{l,n}^{\rm BEM}\right|^{2}\right]$ $=\mathbb{E}\left[\sum_{l=0}^{L-1}\left|\sum_{q=-Q}^{Q}I_{l,q}e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\left(h_{l,q,n}-h_{l,q,n}^{\rm BEM}\right)\right|^{2}\right]$
@@
$\mathbb{E}\left[\sum_{l=0}^{L-1}\left|h_{l,n}-h_{l,n}^{\rm BEM}\right|^{2}\right]<\sum_{l=0}^{L-1}\sum_{q=-Q}^{Q}\mathbb{E}\left[I_{l,q}\right]\frac{N_{\rm D}\,\sigma_{\alpha}^{2}\,A_{W}\epsilon}{2}\leq\epsilon$ (58)
where the second inequality is due to the fact that $A_{W}\leq 2$ and that $\sum_{l=0}^{L-1}\sum_{q=-Q}^{Q}\mathbb{E}\left[I_{l,q}\right]=\frac{1}{N_{\rm D}\,\sigma_{\alpha}^{2}}$ due to the power normalization condition in (18). This completes the proof of the theorem.
@@ Appendix D Proof of Theorem 4
In what follows we use the notation $h_{l,q,n}^{\rm ext}\left(\hat{\mathbf{h}}_{l,q}^{\rm BEM}\right)$ to designate the DPSS extrapolation predictor defined in (37) to highlight its dependence on the estimated vector $\hat{\mathbf{h}}_{l,q}^{\rm BEM}$. When $\hat{\mathbf{h}}_{l,q}^{\rm BEM}$ in (37) is replaced with an arbitrary channel vector $\mathbf{h}$, the DPSS predictor generalizes to
@@
For instance, $h_{l,q,n}^{\rm ext}\left(\mathbf{h}_{l,q}^{\rm BEM}\right)$ is the DPSS predictor based on the actual, not the estimated, vector $\mathbf{h}_{l,q}^{\rm BEM}$. Next, we establish the link between the DPSS extrapolation predictor and MMSE prediction given the knowledge of the channel during the observation interval.
For that sake first note that for each $l$, $q$ and any $n\in\mathbb{Z}$, the random variable $h_{l,q,n}$ as defined by (22) follows a complex symmetric Gaussian distribution $\mathcal{CN}\left(0,N_{\rm D}\sigma_{\alpha}^{2}\right)$ under the conditions of Definition 4. Moreover, the random process $\left(h_{l,q,n}\right)_{n\in\mathbf{Z}}$ is stationary and has an auto-correlation $\mathbb{E}\left[h_{l,q,n}h_{l,q,m}^{*}\right]=\sigma_{\alpha}^{2}N_{\rm D}e^{\imath 2\pi\frac{(n-m)q}{N}}\frac{N}{\pi(n-m)}\sin{(\frac{\pi(n-m)}{N})}$ due to (22). Similarly, $\mathbf{h}_{l,q}\sim\mathcal{CN}\left(\mathbf{0},\sigma_{\alpha}^{2}N_{\rm D}N\mathbf{E}_{\frac{q}{N}}\boldsymbol{\Sigma}\mathbf{E}_{\frac{q}{N}}^{\rm H}\right)$ and $\mathbf{h}_{l,q}^{\rm BEM}\sim\mathcal{CN}\left(\mathbf{0},\sigma_{\alpha}^{2}N_{\rm D}N\mathbf{E}_{\frac{q}{N}}\mathbf{P}^{\rm BEM}\boldsymbol{\Sigma}\mathbf{P}^{\rm BEM}\mathbf{E}_{\frac{q}{N}}^{\rm H}\right)$ with
$\boldsymbol{\Sigma}\triangleq\left[\frac{1}{\pi(n-m)}\sin{\frac{\pi(n-m)}{N}}\right]_{\begin{subarray}{c}n=0\ldots N-1\\ m=0\ldots N-1\end{subarray}}=\left[C_{n,m}^{(N,W)}\right]_{\begin{subarray}{c}n=0\ldots N-1\\ m=0\ldots N-1\end{subarray}}$ (60)
@@
Here $\boldsymbol{\rho}\triangleq\left[e^{-\imath 2\pi\frac{mq}{N}}C_{n,m}^{(N,W)}\right]_{m=0}^{N-1}$. It follows that the reduced-rank MMSE predictor $\hat{h}_{l,q,n}^{\rm RR}$ of rank $Q$ of $h_{l,q,n}$ given $\mathbf{h}_{l,q}$ (where notation ‘$\mathrm{RR}$’ stands for “reduced rank”) is [34]
@@
$\mathbf{f}_{Q}^{\rm RR}\triangleq\mathbf{E}_{\frac{q}{N}}\mathbf{U}_{Q}\mathrm{diag}\left(\frac{1}{\lambda_{1}^{(N,W)}},\ldots,\frac{1}{\lambda_{Q}^{(N,W)}}\right)\mathbf{U}_{Q}^{\rm H}\mathbf{E}_{\frac{q}{N}}^{\rm H}\boldsymbol{\rho}^{\rm H}.$ (63)
@@
Since $\lim_{\sigma_{w}^{2}\to 0}\mathbb{E}\left[\left\|\hat{\mathbf{h}}_{l,q}^{\rm BEM}-\mathbf{h}_{l,q}^{\rm BEM}\right\|^{2}\right]$ due to Assumption 3, it follows from (59) by standard MSE derivations that $\lim_{\sigma_{w}^{2}\to 0}E_{1}=0$. As for $E_{2}$, and since $\mathbf{h}_{l,q}^{\rm BEM}=\mathbf{E}_{\frac{q}{N}}\mathbf{P}^{\rm BEM}\mathbf{E}_{\frac{q}{N}}^{\rm H}\mathbf{h}_{l,q}$ per (25), we have that $\mathbf{U}_{Q_{\rm BEM}}^{\rm H}\mathbf{E}_{\frac{q}{N}}^{\rm H}\mathbf{h}_{l,q}=\mathbf{U}_{Q_{\rm BEM}}^{\rm H}\mathbf{E}_{\frac{q}{N}}^{\rm H}\mathbf{h}_{l,q}^{\rm BEM}$. It follows from (59) that $h_{l,q,n}^{\rm ext}\left(\mathbf{h}_{l,q}\right)=h_{l,q,n}^{\rm ext}\left(\mathbf{h}_{l,q}^{\rm BEM}\right)$ and hence that $E_{2}=0$. Finally, note by referring to (35), (59) and (63) that $\mathbf{f}_{Q}^{\rm RR}=\mathbf{f}_{Q_{\rm BEM}}^{\rm ext}$ if $Q=Q_{\rm BEM}$ leading to $E_{3}=0$.
Now, due to (35), $h_{l,n}^{\rm ext}=\sum_{l=0}^{L-1}I_{l,q}h_{l,q,n}^{\rm ext}$. Moreover, due to the independence conditions from Definition 4, $\hat{h}_{l,n}^{\rm RR}\triangleq\sum_{l=0}^{L-1}I_{l,q}\hat{h}_{l,q,n}^{\rm RR}$ is the reduced-rank MMSE estimate of $h_{l,n}$ given $\left\{\mathbf{h}_{l,q}\right\}_{q=-Q\cdots Q}$ and conditioned on a given realization of $I_{l,q}$. Putting all these pieces together, it follows that $\lim_{\sigma_{w}^{2}\to 0}\left|h_{l,n}^{\rm ext}-\hat{h}_{l,n}^{\rm RR}\right|^{2}$. This completes the proof of the theorem.

### v2

@@ Appendix B Proof of Theorem 2
where $\widetilde{\mathbf{y}}_{\rm p}=[\widetilde{\mathbf{y}}_{{\rm p},0}^{\rm T}\quad\cdots\quad\widetilde{\mathbf{y}}_{{\rm p},(L-1)P_{\rm afdm}}^{\rm T}]^{\rm T}$. For each $l$, $\widetilde{\mathbf{y}}_{{\rm p},l}$ is a $N_{\rm p}\times 1$ vector composed of the pilot samples received at DAFT domain positions $\left\{m_{p}+l\right\}_{p=1\cdots N_{\rm p}}$. Note that by this definition $\tilde{\mathbf{y}}_{\rm p}$ is obtained by permuting $\mathbf{y}_{\rm p}$ in (39) in accordance with the permutation that gives $\tilde{\boldsymbol{\alpha}}$ from $\boldsymbol{\alpha}$. Next, by setting for each $p\in\llbracket 1\mkern 1.5mu..\mkern 1.5muN_{\rm p}\rrbracket$ $m_{p}=q_{p}\frac{N}{2\lceil\frac{Q}{P_{\rm afdm}}\rceil+1}$ for some integer $q_{p}$ it follows from (17) and (55) that $\widetilde{\mathbf{M}}_{\rm p}$ has the following Kronecker structure
$\widetilde{\mathbf{M}}_{\rm p}=\mathbf{I}_{(L-1)P_{\rm afdm}+1}\otimes\widetilde{\mathbf{M}}_{\cal D},$ (58)
$\widetilde{\mathbf{M}}_{\mathcal{D}}=\mathrm{diag}({\rm p}_{1}\cdots{\rm p}_{N_{\rm p}})\mathbf{F}_{2\lceil\frac{Q}{P_{\rm afdm}}\rceil+1,N_{\rm p}}\boldsymbol{\Psi}\>,$ (59)
where $\mathbf{F}_{2\lceil\frac{Q}{P_{\rm afdm}}\rceil+1,N_{\rm p}}$ is a $\left(2\lceil\frac{Q}{P_{\rm afdm}}\rceil+1\right)\times N_{\rm p}$ partial Fourier measurement matrix, and $\boldsymbol{\Psi}$ is a diagonal matrix with unit-modulus entries. We can thus use [37, Theorem 4.5] pertaining to sub-sampled Fourier matrices to conclude that, for sufficiently large $L$, $Q$, sufficiently small $\delta$, and
$N_{\rm p}>\Omega(\frac{1}{\delta^{2}}\log^{2}\frac{1}{\delta}\log\frac{\log(LP_{\rm afdm})}{\delta}\log(LP_{\rm afdm})\log\frac{Q}{P_{\rm afdm}})$ (60)
the RIP constant $\delta_{\tilde{s}_{\rm D}}$ of $\widetilde{\mathbf{M}}_{\mathcal{D}}$ satisfies $\delta_{\tilde{s}_{\rm D}}\leq\delta$ with probability $1-e^{-\Omega\left(\log{\frac{Q}{P_{\rm afdm}}}\log{\frac{1}{\delta}}\right)}$. The RIP of $\mathbf{I}_{(L-1)P_{\rm afdm}+1}$ trivially satisfies $\delta_{\tilde{s}_{\rm d}}=0$. As for the HiRIP of $\widetilde{\mathbf{M}}_{\rm p}$, we can apply [24, Theorem 4] to (58) thanks to its Kronecker structure to conclude that, if $N_{\rm p}$ and $\delta$ are as in (60), then
@@ Appendix C Proof of Theorem 3
We use the following lemma taken from [38, Theorem 2.4].
@@
[38, Theorem 2.4] Let $h(t)$ be a continuous-time zero-mean wide-sense stationary random process with power spectrum $P_{h}(f)=\frac{1}{B}\mathrm{rect}(\frac{f-F_{\rm c}}{2B})$. Denote by $\mathbf{h}=[h(0T_{\rm s})\,\ldots\>h((N-1)T_{s})]^{\rm T}$ a vector of samples acquired from $h(t)$ with a sampling period $T_{s}\leq\frac{1}{2F_{\rm c}+B}$. Let $W=\frac{BT_{s}}{2}$, $\mathbf{U}_{k}$ the matrix form of the $k$ first $(N,W)$-DPSS vectors, $\mathbf{E}_{f}\triangleq\mathrm{diag}\left(e^{\imath 2\pi f0}\ldots e^{\imath 2\pi f(N-1)}\right)$ and $\mathbf{P}_{k}\triangleq\mathbf{E}_{F_{\rm c}T_{\rm s}}\mathbf{U}_{k}\mathbf{U}_{k}^{\rm H}\mathbf{E}_{F_{\rm c}T_{\rm s}}^{\rm H}$. Then $\mathbb{E}[\left\|\mathbf{h}-\mathbf{P}_{k}\mathbf{h}\right\|_{2}^{2}]=\frac{1}{2W}\sum_{l=k}^{N-1}\lambda_{l}^{(N,W)}$.
In what follows, we use Lemma 2 to upper bound $\mathbb{E}[|h_{l,q,n}-h_{l,q,n}^{\rm BEM}|^{2}]$ by first rewriting $h_{l,q,n}$ as the sampled version of the continuous-time signal $h_{l,q}(t)$ defined as
@@
with $f_{i}\triangleq\frac{q}{NT_{\rm s}}+\frac{\kappa_{i}}{NT_{\rm s}}$. To prove that the PSD of the random process $h_{l,q}(t)$ has the desired property, we derive its autocorrelation function $R_{h_{l,q}}(\tau)\triangleq\mathbb{E}[h_{l,q}(t)h_{l,q}^{*}(t+\tau)]$
$R_{h_{l,q}}(\tau)$ $=\sum_{i=1}^{N_{D}}\sum_{j=1}^{N_{D}}\mathbb{E}[\alpha_{l,q,i}\alpha_{l,q,j}^{*}]\,\mathbb{E}[e^{\imath 2\pi f_{i}t}e^{-\imath 2\pi f_{j}(t+\tau)}]$
@@
where the first equality follows from (62) and the second from $f_{i}\sim\mathcal{U}([\frac{q}{NT_{\rm s}}-\frac{1}{2NT_{\rm s}},\frac{q}{NT_{\rm s}}+\frac{1}{2NT_{\rm s}}])$, $\alpha_{l,q,i}\sim\mathcal{CN}(0,\sigma_{\alpha}^{2})$ and the independence of $\left\{\alpha_{l,q,i}\right\}_{i}$ as per Definition 4. This gives:
@@
Power spectral density (PSD) $P_{h_{l,q}}(f)\triangleq\mathcal{F}\{R_{h_{l,q}}(\tau)\}$ is thus
$P_{h_{l,q}}(f)=N_{D}\,\sigma_{\alpha}^{2}\,NT_{\rm s}\,\text{rect}\left(\left(f-\frac{q}{NT_{\rm s}}\right)NT_{\rm s}\right).$ (65)
@@
[…]
Now define $\lambda_{b}^{(c)}$ as the $b$-th eigenvalue of the prolate spheroidal wave functions (PSWF) [39] with the bandwidth parameter $c\triangleq\pi NW$. This allows us to exploit existing results on the behavior of PSWF eigenvalues in the limit of $c\to\frac{\pi}{2}$, i.e., as $W\to 0$ at rate $\frac{1}{2N}$, to upper bound the sum of DPSS eigenvalues $\lambda_{b}^{(N,W)}$ in (66). Indeed, due to [40, Theorem 2]
@@
Here, $A_{W}$ is a function of $W$ defined in [40, Eq. (45)] and its image is fully included in the interval $\frac{\pi^{2}}{8},2]$. Plugging (67) into (66) and noting that $2WN=1$ leads to
@@
The right-hand side term in (68) can be upper bounded due to the fact that the PSWF eigenvalues decay at least exponentially  [^2] Actually, even super-geometric decay can be proven [41] as $b$ grows beyond $\frac{2\pi NW}{\pi}+O(\log(\pi NW))=1+O(\log\frac{\pi}{2})$. More precisely, it follows from [42, Theorem 2.5] that $\lambda_{b}^{(c)}=O(e^{-\frac{\pi^{2}}{\log\frac{\pi}{2}}b})$. Plugging this into (68) gives, for any $\epsilon>0$ and $Q_{\rm BEM}>C\log\frac{2}{\epsilon}$ for sufficiently large $C$,
@@
$\mathbb{E}\left[\sum_{l=0}^{L-1}\left|h_{l,n}-h_{l,n}^{\rm BEM}\right|^{2}\right]$
$=\mathbb{E}\left[\sum_{l=0}^{L-1}\left|\sum_{q=-Q}^{Q}I_{l,q}e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\left(h_{l,q,n}-h_{l,q,n}^{\rm BEM}\right)\right|^{2}\right]$
@@
$\mathbb{E}\left[\sum_{l=0}^{L-1}\left|h_{l,n}-h_{l,n}^{\rm BEM}\right|^{2}\right]<\sum_{l,q}\mathbb{E}\left[I_{l,q}\right]\frac{N_{\rm D}\,\sigma_{\alpha}^{2}\,A_{W}\epsilon}{2}\leq\epsilon$ (71)
where the second inequality is due to the fact that $A_{W}\leq 2$ and that $\sum_{l=0}^{L-1}\sum_{q=-Q}^{Q}\mathbb{E}\left[I_{l,q}\right]=\frac{1}{N_{\rm D}\,\sigma_{\alpha}^{2}}$ is due to the power normalization condition in (28). This completes the proof.
@@ Appendix D Proof of Theorem 4
We now use the notation $h_{l,q,n}^{\rm ext}\left(\hat{\mathbf{h}}_{l,q}^{\rm BEM}\right)$ for the DPSS extrapolation predictor defined in (47) to highlight its dependence on the estimate $\hat{\mathbf{h}}_{l,q}^{\rm BEM}$. When $\hat{\mathbf{h}}_{l,q}^{\rm BEM}$ in (47) is replaced with an arbitrary vector $\mathbf{h}$, the DPSS predictor generalizes to
@@
For instance, $h_{l,q,n}^{\rm ext}\left(\mathbf{h}_{l,q}^{\rm BEM}\right)$ is the DPSS predictor based on the actual, not the estimated, vector $\mathbf{h}_{l,q}^{\rm BEM}$. Next, we establish the link between the DPSS extrapolation predictor and an MMSE predictor that uses the knowledge of the channel values during the observation interval. For that sake, first note that for each $l$, $q$ and any $n\in\mathbb{Z}$, the random variable $h_{l,q,n}$ as defined by (32) follows a complex symmetric Gaussian distribution $\mathcal{CN}\left(0,N_{\rm D}\sigma_{\alpha}^{2}\right)$ under the conditions of Definition 4. Moreover, the random process $\left(h_{l,q,n}\right)_{n\in\mathbf{Z}}$ is stationary and has an auto-correlation $\mathbb{E}[h_{l,q,n}h_{l,q,m}^{*}]=\sigma_{\alpha}^{2}N_{\rm D}e^{\imath 2\pi\frac{(n-m)q}{N}}\frac{N}{\pi(n-m)}\sin{(\frac{\pi(n-m)}{N})}$ due to (32). Similarly, $\mathbf{h}_{l,q}\sim\mathcal{CN}(\mathbf{0},\sigma_{\alpha}^{2}N_{\rm D}N\mathbf{E}_{\frac{q}{N}}\boldsymbol{\Sigma}\mathbf{E}_{\frac{q}{N}}^{\rm H})$ and $\mathbf{h}_{l,q}^{\rm BEM}\sim\mathcal{CN}(\mathbf{0},\sigma_{\alpha}^{2}N_{\rm D}N\mathbf{E}_{\frac{q}{N}}\mathbf{P}^{\rm BEM}\boldsymbol{\Sigma}\mathbf{P}^{\rm BEM}\mathbf{E}_{\frac{q}{N}}^{\rm H})$ with
$\boldsymbol{\Sigma}\triangleq\Bigg[\underbrace{\frac{1}{\pi(n-m)}\sin{\frac{\pi(n-m)}{N}}}_{C_{n,m}^{(N,W)}}\Bigg]_{\begin{subarray}{c}n=0\ldots N-1\\ m=0\ldots N-1\end{subarray}}$ (73)
@@
Here, $\boldsymbol{\rho}\triangleq[e^{-\imath 2\pi\frac{mq}{N}}C_{n,m}^{(N,W)}]_{m=0}^{N-1}$. It follows that the reduced-rank MMSE predictor $\hat{h}_{l,q,n}^{\rm RR}$ of rank $Q$ of $h_{l,q,n}$ given $\mathbf{h}_{l,q}$ (where notation $\mathrm{RR}$’ stands for “reduced rank”) is [43]
@@
$\mathbf{f}_{Q}^{\rm RR}\triangleq\mathbf{E}_{\frac{q}{N}}\mathbf{U}_{Q}\mathrm{diag}(\frac{1}{\lambda_{1}^{(N,W)}},\ldots,\frac{1}{\lambda_{Q}^{(N,W)}})\mathbf{U}_{Q}^{\rm H}\mathbf{E}_{\frac{q}{N}}^{\rm H}\boldsymbol{\rho}^{\rm H}.$ (76)
@@
Since $\lim_{\sigma_{w}^{2}\to 0}\mathbb{E}[\|\hat{\mathbf{h}}_{l,q}^{\rm BEM}-\mathbf{h}_{l,q}^{\rm BEM}\|^{2}]$ due to Assumption 3, it follows from (72) by standard MSE derivations that $\lim_{\sigma_{w}^{2}\to 0}E_{1}=0$. As for $E_{2}$, and since $\mathbf{h}_{l,q}^{\rm BEM}=\mathbf{E}_{\frac{q}{N}}\mathbf{P}^{\rm BEM}\mathbf{E}_{\frac{q}{N}}^{\rm H}\mathbf{h}_{l,q}$ per (35), we have that $\mathbf{U}_{Q_{\rm BEM}}^{\rm H}\mathbf{E}_{\frac{q}{N}}^{\rm H}\mathbf{h}_{l,q}=\mathbf{U}_{Q_{\rm BEM}}^{\rm H}\mathbf{E}_{\frac{q}{N}}^{\rm H}\mathbf{h}_{l,q}^{\rm BEM}$. It follows from (72) that $h_{l,q,n}^{\rm ext}\left(\mathbf{h}_{l,q}\right)=h_{l,q,n}^{\rm ext}\left(\mathbf{h}_{l,q}^{\rm BEM}\right)$ and hence that $E_{2}=0$. Finally, note by referring to (45), (72) and (76) that $\mathbf{f}_{Q}^{\rm RR}=\mathbf{f}_{Q_{\rm BEM}}^{\rm ext}$ if $Q=Q_{\rm BEM}$ leading to $E_{3}=0$. Now, due to (45), $h_{l,n}^{\rm ext}=\sum_{l=0}^{L-1}I_{l,q}h_{l,q,n}^{\rm ext}$. Moreover, due to the independence conditions from Definition 4, $\hat{h}_{l,n}^{\rm RR}\triangleq\sum_{l=0}^{L-1}I_{l,q}\hat{h}_{l,q,n}^{\rm RR}$ is the reduced-rank MMSE estimate of $h_{l,n}$ given $\left\{\mathbf{h}_{l,q}\right\}_{q=-Q\cdots Q}$ and conditioned on a given realization of $I_{l,q}$. Putting all these pieces together, it follows that $\lim_{\sigma_{w}^{2}\to 0}|h_{l,n}^{\rm ext}-\hat{h}_{l,n}^{\rm RR}|^{2}$. This completes the proof.